The solubility of Al(OH)3 is 5.15 x 10-7 g/ℓ. Calculate the Ksp of Al(OH)3.


1
Expert's answer
2021-08-16T02:46:02-0400

Solution:

Convert g/L to mol/L:

The molar mass of Al(OH)3 is 78 g/mol.

Therefore,

(5.15×10-7 g Al(OH)3 / 1 L) × (1 mol Al(OH)3 / 78 g Al(OH)3) = 6.6×10-9 mol/L

S(Al(OH)3) = 6.6×10-9 mol/L


Al(OH)3(s) ⇌ Al3+(aq) + 3OH-(aq)

According to the equation above:

[Al3+] = S(Al(OH)3) = 6.6×10-9 mol/L

[OH-] = 3 × S(Al(OH)3) = 3 × (6.6×10-9 mol/L) = 1.98×10-8 mol/L


The Ksp expression for Al(OH)3(s) is:

Ksp = [Al3+] × [OH-]3

Ksp = (6.6×10-9) × (1.98×10-8)3 = 5.12×10-32

Ksp = 5.12×10-32


Answer: Ksp of Al(OH)3 is 5.12×10-32